# Nonlocal pseudo-symmetries and Bäcklund transformations as $\mathcal{C}$-morphisms

### v2

@@ 4.1 Examples
In this subsection, we will analyze several examples of equations with two independent variables for which, starting from a ZCR, it is possible to determine pseudosymmetries and subsequently ABT and BT, following the procedure described above. Moreover, in the last example, we will use a $2$-pseudosymmetry to determine an ABT of the Tzitzeica equation.
@@
We consider here KdV =$\mathcal{E}=\{z_{xxx}=-z_{t}-6zz_{x}\}$ and show how the invariants of a nonlocal pseudosymmetry $Y$ can be used to determine an ABT of this equation. As already shown in Example 24, $\mathcal{E}$ admits the nonlocal conservation law $\mu=\rho dx+\left(z_{x}+\left(4\lambda-2z\right)\rho\right)dt$ in the $1$-dimensional differentiable covering $\mathcal{Y}^{(\infty)}\rightarrow\mathcal{E}^{(\infty)}$ defined by
$\mathcal{Y}:\;\left\{\begin{array}[]{l}\rho_{x}=-\rho^{2}-z-\lambda,\vspace{5pt}\\ \rho_{t}=-2(2\lambda-z)\rho^{2}-2z_{x}\rho+z_{xx}+2z^{2}-2\lambda z-4\lambda^{2},\vspace{5pt}\\ z_{xxx}=-z_{t}-6zz_{x}.\end{array}\right.$
@@
First, by searching for pseudosymmetries of $\mathcal{Y}$ that are pseudo-prolongations of a vector field
@@
relatively to conservation laws $c\,\mu$, with $c\in\mathbb{R}$, one finds that the most general pseudosymmetry of this type is a constant multiple of
$Y=\rho\partial_{z}-\frac{1}{4}\partial_{\rho}+\;\ldots,$
with $c=2$. Such a pseudosymmetry of $\mathcal{Y}$ can be seen as a nonlocal pseudosymmetry of $\mathcal{E}$, with respect to the differentiable covering defined by $\mathcal{Y}$. Thus, being $\{x,\;t,\;z+2\rho^{2}\}$ a basic system of zeroth order invariants of $Y$, one has that the ABT of KdV discussed in Example 18
@@
is of the type described above, i.e., it is provided by the invariants of above nonlocal pseudosymmetry $Y$ of $\mathcal{E}$.
@@
We show here that also the Darboux transformation discussed in Example 20 is provided by the invariants of a nonlocal pseudosymmetry. To this end one can consider the $2$-dimensional differentiable covering of KdV =$\mathcal{E}=\{z_{xxx}=-z_{t}-6zz_{x}\}$ defined by
$\mathcal{Y}:\;\left\{\begin{array}[]{l}\hat{\rho}_{x}=-\hat{\rho}^{2}-z-\hat{\lambda},\vspace{5pt}\\ \hat{\rho}_{t}=-2(2\hat{\lambda}-z)\hat{\rho}^{2}-2z_{x}\hat{\rho}+z_{xx}+2z^{2}-2\hat{\lambda}z-4\hat{\lambda}^{2},\vspace{5pt}\\ \rho_{x}=-\rho^{2}-z-\lambda,\vspace{5pt}\\ \rho_{t}=-2(2\lambda-z)\rho^{2}-2z_{x}\rho+z_{xx}+2z^{2}-2\lambda z-4\lambda^{2},\vspace{5pt}\\ z_{xxx}=-z_{t}-6zz_{x},\end{array}\right.$
which coincides with (28) and is obtained by doubling the Riccati system of previous example. Hence, being $\mu=\rho dx+\left(z_{x}+\left(4\lambda-2z\right)\rho\right)dt$ still a conservation law for $\mathcal{Y}$, one can search for pseudosymmetries of $\mathcal{Y}$ that are pseudo-prolongations of a vector field
@@
relatively to $c\,\mu$, with $c\in\mathbb{R}$. In this case, when $\hat{\lambda}\neq\lambda$ and $\hat{\rho}\neq\rho$, one finds that the most general pseudosymmetry of this type is a constant multiple of
$Y=\hat{\rho}\partial_{z}-\frac{1}{4}\left(1+\frac{(\rho-\hat{\rho})^{2}}{\lambda-\hat{\lambda}}\right)\partial_{\rho}-\frac{1}{4}\partial_{\hat{\rho}}+...,$
with $c=2$. Such a pseudosymmetry of $\mathcal{Y}$ can be seen as a nonlocal pseudosymmetry of $\mathcal{E}$, with respect to the differentiable covering defined by $\mathcal{Y}$. Thus, being $\{x,\;t,\;z+2\hat{\rho}^{2},\hat{\rho}+(\hat{\lambda}-\lambda)/(\hat{\rho}-\rho)\,\}$ a basic system of zeroth order invariants of $Y$, one has that also the Darboux transformation discussed in Example 20
$x^{\prime}=x,\qquad t^{\prime}=t,\qquad z^{\prime}=-z-2\hat{\rho}^{2}{\footnotesize-2\hat{\lambda}},\qquad\rho^{\prime}=-\frac{\hat{\lambda}-\lambda}{\hat{\rho}-\rho}-\hat{\rho}$
is provided by the invariants of a nonlocal pseudosymmetry of $\mathcal{E}$. In this case the transformation describes a regular $\mathcal{C}$-morphism from $\mathcal{Y}$ to $\mathcal{E}^{\prime}$, where by $\mathcal{E}^{\prime}$ now we mean the system
:$\mathcal{E}^{\prime}:\;\left\{\begin{array}[]{l}\rho^{\prime}_{x}=-\rho^{\prime}{}^{2}-z^{\prime}-\lambda,\vspace{5pt}\\ \rho^{\prime}_{t}=-2(2\lambda-z^{\prime})\rho^{\prime}{}^{2}-2z^{\prime}_{x}\rho^{\prime}+z^{\prime}_{xx}+2z^{\prime}{}^{2}-2\lambda z^{\prime}-4\lambda^{2},\vspace{5pt}\\ z^{\prime}_{xxx}=-6z^{\prime}z^{\prime}_{x}-z^{\prime}_{t}.\end{array}\right.$
@@
We consider here sine-Gordon equation =$\mathcal{E}=\{z_{xt}=\sin z\}$ and show how the invariants of a nonlocal pseudosymmetry $Y$ can be used to determine an ABT of this equation. To this end we consider the following ZCR
$\alpha=\frac{1}{2}\left(\begin{array}[]{ll}\eta&-z_{x}\vspace{5pt}\\ z_{x}&-{\displaystyle\eta}\end{array}\right)dx+\frac{1}{2\eta}\left(\begin{array}[]{ll}\cos\,z\vspace{5pt}&\sin\,z\\ \sin\,z&-\cos\,z\end{array}\right)dt,\qquad\eta\in\mathbb{R}-\{0\}.$
@@
$\mathcal{Y}:\;\left\{\begin{array}[]{l}\rho_{x}=-\frac{1}{2}z_{x}\rho^{2}+\eta\rho-\frac{1}{2}\,z_{x}\vspace{5pt}\\ \rho_{t}=\frac{1}{2\eta}\sin z\,\left(1-\rho^{2}\right)+\frac{1}{\eta}\cos z\,\rho\vspace{5pt}\\ z_{xt}=\sin z\end{array}\right.$
@@
First, by searching for pseudosymmetries of $\mathcal{Y}$ that are pseudo-prolongations of a vector field
@@
relatively to conservation laws $c\,\mu$, with $c\in\mathbb{R}$, one finds that the most general pseudosymmetry of this type is a constant multiple of
$Y=\left(\rho^{2}+1\right)\partial_{z}+\frac{\rho^{2}+1}{4}\,\partial_{\rho}+\;\ldots,$
with $c=2$. Such a pseudosymmetry of $\mathcal{Y}$ can be seen as a nonlocal pseudosymmetry of $\mathcal{E}$, with respect to the differentiable covering defined by $\mathcal{Y}$. Thus, being $\{x,\,t,\,z+4\,\arctan\rho\}$ a basic system of zeroth order invariants of $Y$, one is naturally lead to consider $\xi^{1}=x,\,\xi^{2}=t$ and $\nu=\nu\left(z+4\,\arctan\rho\right)$. This way, a direct computation shows that $z^{\prime}_{xt}-\sin z^{\prime}=0\mod{\mathcal{E}^{(\infty)}}$ iff $\nu=z+4\,\arctan\rho$. Hence, one has the ABT of $\mathcal{E}$
@@
provided by the invariants of above nonlocal pseudosymmetry $Y$ of $\mathcal{E}$.
@@
We consider here the short-pulse equation =$\mathcal{E}=\{z_{xt}=\frac{1}{2}z^{2}z_{xx}+zz_{x}^{2}+z\}$ and show how the invariants of a nonlocal pseudosymmetry $Y$ can be used to determine an ABT of this equation. To this end we consider the following ZCR
$\alpha=\frac{1}{4\eta}\left(\begin{array}[]{ll}1&z_{x}\vspace{5pt}\\ z_{x}&-1\end{array}\right)dx+\frac{1}{8\eta}\left(\begin{array}[]{ll}8\eta^{2}+z^{2}&z\left(zz_{x}-4\,\eta\right)\vspace{5pt}\\ z\left(zz_{x}+4\,\eta\right)&-8\eta^{2}-z^{2}\end{array}\right)dt,\qquad\eta\in\mathbb{R}-\{0\}.$
@@
$\mathcal{Y}:\;\left\{\begin{array}[]{l}{\displaystyle\rho_{x}=-\frac{z_{x}}{4\eta}\rho^{2}+\frac{1}{2\eta}\,\rho+\frac{z_{x}}{4\eta}},\vspace{10pt}\\ {\displaystyle\rho_{t}=-\frac{z\left(zz_{x}+4\,\eta\right)}{8\eta}\rho^{2}+\frac{\left(8\,\eta^{2}+z^{2}\right)}{4\eta}\rho-\frac{z\left(-zz_{x}+4\,\eta\right)}{8\eta}},\vspace{10pt}\\ {\displaystyle z_{xt}=\frac{1}{2}z^{2}z_{xx}+zz_{x}^{2}+z,}\end{array}\right.$
@@
First, by searching for pseudosymmetries of $\mathcal{Y}$ that are pseudo-prolongations of a vector field
@@
relatively to conservation laws $c\,\mu$, with $c\in\mathbb{R}$, one finds that the most general pseudosymmetry of this type is a constant multiple of
$Y=-2\rho\partial_{x}+\left(\rho^{2}-1\right)\partial_{z}-\frac{1}{8\eta}\left(\rho^{2}+1\right)^{2}\partial_{\rho}+...,$
with $c=2$. Such a pseudosymmetry of $\mathcal{Y}$ can be seen as a nonlocal pseudosymmetry of $\mathcal{E}$, with respect to the differentiable covering defined by $\mathcal{Y}$. Thus, being $\{x+8\eta/(\rho^{2}+1),\,t,z-8\eta\rho/(\rho^{2}+1)\}$ a basic system of zeroth order invariants of $Y$, one is naturally lead to consider $\xi^{1}=x+8\eta/(\rho^{2}+1),\,\xi^{2}=t$ and $\nu=\nu\left(z-8\eta\rho/(\rho^{2}+1)\right)$. This way, a direct computation shows that $z^{\prime}_{x^{\prime}t^{\prime}}-\frac{1}{2}z^{\prime}{}^{2}z^{\prime}_{x^{\prime}x^{\prime}}-z^{\prime}z^{\prime}{}_{x^{\prime}}^{2}-z^{\prime}=0\mod{\mathcal{E}^{(\infty)}}$ iff $\nu=-z+8\eta\rho/(\rho^{2}+1)$. Hence, one has the ABT of $\mathcal{E}$
@@
provided by the invariants of above nonlocal pseudosymmetry $Y$ of $\mathcal{E}$.
@@
We consider here Camassa-Holm equation =$\mathcal{E}=\{z_{txx}=3\,z\,z_{x}-z\,z_{xxx}-2\,z_{x}\,z_{xx}+z_{t}\}$ and show how the invariants of a nonlocal pseudosymmetry $Y$ can be used to determine an ABT of this equation. To this end we consider the following ZCR
$\begin{array}[]{ll}\alpha={\displaystyle\frac{1}{2}}\left(\begin{array}[]{ll}0\vspace{5pt}&-{\displaystyle\frac{1}{\eta}}\\ -z+z_{xx}-\eta&\quad 0\end{array}\right)\,dx+{\displaystyle\frac{1}{2}}\left(\begin{array}[]{ll}z_{x}\vspace{5pt}&{\displaystyle\frac{1}{\eta}}(z-2\,\eta)\\ z^{2}-z\,z_{xx}-\eta\,z-2\eta^{2}&\quad-z_{x}\end{array}\right)dt,&{\scriptsize\eta\in\mathbb{R}-\{0\}.}\end{array}$
@@
$\mathcal{Y}:\;\left\{\begin{array}[]{l}\rho_{x}={\displaystyle\frac{1}{2\eta}\rho^{2}+\frac{z_{xx}}{2}+\frac{-{\eta}^{2}-\eta\,z}{2\,\eta}},\vspace{10pt}\\ \rho_{t}={\displaystyle-\frac{\left(-2\,\eta+z\right)}{2\,\eta}\rho^{2}-z_{x}\,\rho-\frac{z\,z_{xx}}{2}-\frac{2\,\eta^{3}+\eta^{2}z-\eta\,z^{2}}{2\,\eta}},\vspace{10pt}\\ z_{txx}=3\,z\,z_{x}-z\,z_{xxx}-2\,z_{x}\,z_{xx}+z_{t},\end{array}\right.$ (35)
@@
In this case, if one searches for pseudosymmetries of $\mathcal{Y}$ that are pseudo-prolongations of a vector field
@@
relatively to conservation laws $c\,\mu$, with $c\in\mathbb{R}$, one finds that the most general pseudosymmetry of this type is a constant multiple of
$Y=-\partial_{x}+(2\rho-z_{x})\,\partial_{z}+\frac{\left(\eta^{2}-\rho^{2}\right)}{2\eta}\partial_{\rho}+...\,,$
@@
$\xi^{1}=x-\ln\left(\frac{\rho-\eta}{\rho+\eta}\right),\quad\xi^{2}=t,\quad\nu^{1}=\left(\frac{\rho^{2}+\eta^{2}}{\rho^{2}-\eta^{2}}\right)z+\frac{2\,\eta\,\rho}{\rho^{2}-\eta^{2}}\,z_{x}\qquad\nu^{2}=\frac{2\,\eta\,\rho}{\rho^{2}-\eta^{2}}z+\left(\frac{\rho^{2}+\eta^{2}}{\rho^{2}-\eta^{2}}\right)\,z_{x}$
a basic system of first-order order invariants of $Y$, one finds that in this case the system $\mathcal{Q}$ (the factorization of $\mathcal{E}$ by $Y$) reads
$\mathcal{Q}:\quad\left\{\begin{array}[]{l}\nu_{\xi^{1}}^{1}=\nu^{2},\vspace{10pt}\\ \nu_{\xi^{1}\xi^{2}}^{2}=\left(-6\,\eta+3\,\nu^{1}-2\,\nu_{\xi^{1}}^{2}\right)\nu^{2}+\left(2\,\eta-\nu^{1}\right)\nu_{\xi^{1}\xi^{1}}^{2}+\nu_{\xi^{2}}^{1}.\vspace{10pt}\end{array}\right.$
@@
one gets that $z^{\prime}_{t^{\prime}x^{\prime}x^{\prime}}-3\,z^{\prime}\,z^{\prime}_{x^{\prime}}+z^{\prime}\,z^{\prime}_{x^{\prime}x^{\prime}x^{\prime}}+2\,z^{\prime}_{x^{\prime}}\,z^{\prime}_{x^{\prime}x^{\prime}}-z^{\prime}_{t^{\prime}}=0\mod{\mathcal{E}^{(\infty)}}$. Hence, by means of (37), the invariants of nonlocal pseudosymmetry $Y$ of $\mathcal{E}$ provides an ABT of $\mathcal{E}$. This ABT coincides with that found in [36], by means of a different approach.

### v5

@@ 4.1 Examples
In this subsection, we will analyze several examples of equations with two independent variables for which, starting from a ZCR, it is possible to determine pseudosymmetries and subsequently auto-Bäcklund transformations and Bäcklund transformations, following the procedure described above. Moreover, in the last example, we will use a $2$-pseudosymmetry to determine an auto-Bäcklund transformation of the Tzitzeica equation. Throughout this section $x_{1}=x$ and $x_{2}=t$.
@@
We consider here KdV $\mathcal{E}=\{z_{xxx}=-z_{t}-6zz_{x}\}$ and show how the invariants of a nonlocal pseudosymmetry $\bar{Y}$ can be used to determine an auto-Bäcklund transformation of this equation. As already shown in Example 25, $\mathcal{E}$ admits the nonlocal conservation law $\mu=\rho dx+\left(z_{x}+\left(4\lambda-2z\right)\rho\right)dt$ in the $1$-dimensional differentiable covering $\mathcal{Y}^{(\infty)}\rightarrow\mathcal{E}^{(\infty)}$ defined by
$\mathcal{Y}:\;\left\{\begin{array}[]{l}\rho_{x}=-\rho^{2}-z-\lambda,\vskip 5.0pt\\ \rho_{t}=-2(2\lambda-z)\rho^{2}-2z_{x}\rho+z_{xx}+2z^{2}-2\lambda z-4\lambda^{2},\vskip 5.0pt\\ z_{xxx}=-z_{t}-6zz_{x}.\end{array}\right$.
@@
First, by searching for pseudosymmetries $\bar{Y}:=\left.Y\right|_{\mathcal{Y}^{(\infty)}}$ of $\mathcal{Y}$ determined by the pseudoprolongations $Y$ of vector fields of the form
@@
relatively to conservation laws $c\,\mu$, with $c\in\mathbb{R}-\{0\}$, one finds that the most general pseudosymmetry of this type with $b^{1}\neq 0$ is a constant multiple of
$\bar{Y}=\rho\partial_{z}-\frac{1}{4}\partial_{\rho}+\ldots\,$,
with $c=2$. Such a pseudosymmetry of $\mathcal{Y}$ can be seen as a nonlocal pseudosymmetry of $\mathcal{E}$, with respect to the differentiable covering defined by $\mathcal{Y}$. Thus, being $\{x,\;t,\;z+2\rho^{2}\}$ a basic system of zeroth order invariants of $\bar{Y}$, one has that the auto-Bäcklund transformation of KdV discussed in Example 19
@@
is of the type described above, i.e., it is provided by the invariants of the pseudosymmetry $\bar{Y}$.
@@
We show here that also the Darboux transformation discussed in Example 21 is provided by the invariants of a nonlocal pseudosymmetry $\bar{Y}$. To this end one can consider the $2$-dimensional differentiable covering of KdV $\mathcal{E}=\{z_{xxx}=-z_{t}-6zz_{x}\}$ defined by
$\mathcal{Y}:\;\left\{\begin{array}[]{l}\hat{\rho}_{x}=-\hat{\rho}^{2}-z-\hat{\lambda},\vskip 5.0pt\\ \hat{\rho}_{t}=-2(2\hat{\lambda}-z)\hat{\rho}^{2}-2z_{x}\hat{\rho}+z_{xx}+2z^{2}-2\hat{\lambda}z-4\hat{\lambda}^{2},\vskip 5.0pt\\ \rho_{x}=-\rho^{2}-z-\lambda,\vskip 5.0pt\\ \rho_{t}=-2(2\lambda-z)\rho^{2}-2z_{x}\rho+z_{xx}+2z^{2}-2\lambda z-4\lambda^{2},\vskip 5.0pt\\ z_{xxx}=-z_{t}-6zz_{x},\end{array}\right$.
which coincides with (32) and is obtained by doubling the Riccati-type system of previous example. Hence, being $\mu=\rho dx+\left(z_{x}+\left(4\lambda-2z\right)\rho\right)dt$ still a conservation law for $\mathcal{Y}$, one can search for pseudosymmetries $\bar{Y}:=\left.Y\right|_{\mathcal{Y}^{(\infty)}}$ of $\mathcal{Y}$ determined by the pseudoprolongations $Y$ of vector fields of the form
@@
relatively to $2\,\mu$. In this case, when $\hat{\lambda}\neq\lambda$ and $\hat{\rho}\neq\rho$, one finds that the most general pseudosymmetry of this type with $b^{1}\neq 0$ is a constant multiple of
$\bar{Y}=\hat{\rho}\partial_{z}-\frac{1}{4}\left(1+\frac{(\rho-\hat{\rho})^{2}}{\lambda-\hat{\lambda}}\right)\partial_{\rho}-\frac{1}{4}\partial_{\hat{\rho}}$$\ldots\,$.
Such a pseudosymmetry of $\mathcal{Y}$ can be seen as a nonlocal pseudosymmetry of $\mathcal{E}$, with respect to the differentiable covering defined by $\mathcal{Y}$. Thus, being $\{x,\;t,\;z+2\hat{\rho}^{2},\hat{\rho}+(\hat{\lambda}-\lambda)/(\hat{\rho}-\rho)\,\}$ a basic system of zeroth order invariants of $\bar{Y}$, one has that the Darboux transformation discussed in Example 21
$x^{\prime}=x,\qquad t^{\prime}=t,\qquad z^{\prime}=-z-2\hat{\rho}^{2}{\footnotesize-2\hat{\lambda}},\qquad\rho^{\prime}=-\frac{\hat{\lambda}-\lambda}{\hat{\rho}-\rho}-\hat{\rho}$,
is provided by the invariants of the pseudosymmetry $\bar{Y}$. In this case the transformation describes a regular $\mathcal{C}$-morphism from $\mathcal{Y}$ to $\mathcal{E}^{\prime}$, where by $\mathcal{E}^{\prime}$ now we mean the system
$\mathcal{E}^{\prime}:\;\left\{\begin{array}[]{l}\rho^{\prime}_{x}=-\rho^{\prime}{}^{2}-z^{\prime}-\lambda,\vskip 5.0pt\\ \rho^{\prime}_{t}=-2(2\lambda-z^{\prime})\rho^{\prime}{}^{2}-2z^{\prime}_{x}\rho^{\prime}+z^{\prime}_{xx}+2z^{\prime}{}^{2}-2\lambda z^{\prime}-4\lambda^{2},\vskip 5.0pt\\ z^{\prime}_{xxx}=-6z^{\prime}z^{\prime}_{x}-z^{\prime}_{t}.\end{array}\right$.
@@
We consider here sine-Gordon equation $\mathcal{E}=\{z_{xt}=\sin z\}$ and show how the invariants of a nonlocal pseudosymmetry $\bar{Y}$ can be used to determine an auto-Bäcklund transformation of this equation. To this end we consider the following ZCR
$\alpha=\frac{1}{2}\left(\begin{array}[]{ll}\eta&-z_{x}\vskip 5.0pt\\ z_{x}&-{\displaystyle\eta}\end{array}\right)dx+\frac{1}{2\eta}\left(\begin{array}[]{ll}\cos\,z\vskip 5.0pt&\sin\,z\\ \sin\,z&-\cos\,z\end{array}\right)dt,\qquad\eta\in\mathbb{R}-\{0\}.$
@@
$\mathcal{Y}:\;\left\{\begin{array}[]{l}\rho_{x}=-\frac{1}{2}z_{x}\rho^{2}+\eta\rho-\frac{1}{2}\,z_{x},\vskip 5.0pt\\ \rho_{t}=\frac{1}{2\eta}\sin z\,\left(1-\rho^{2}\right)+\frac{1}{\eta}\cos z\,\rho,\vskip 5.0pt\\ z_{xt}=\sin z,\end{array}\right$.
@@
First, by searching for pseudosymmetries $\bar{Y}:=\left.Y\right|_{\mathcal{Y}^{(\infty)}}$ of $\mathcal{Y}$ determined by the pseudoprolongations $Y$ of vector fields of the form
@@
relatively to conservation laws $c\,\mu$, with $c\in\mathbb{R}-\{0\}$, one finds that the most general pseudosymmetry of this type with $b^{1}\neq 0$ is a constant multiple of
$\bar{Y}=\left(\rho^{2}+1\right)\partial_{z}+\frac{\rho^{2}+1}{4}\,\partial_{\rho}+\ldots\,$,
with $c=2$. Such a pseudosymmetry of $\mathcal{Y}$ can be seen as a nonlocal pseudosymmetry of $\mathcal{E}$, with respect to the differentiable covering defined by $\mathcal{Y}$. Thus, being $\{x,\,t,\,z+4\,\arctan\rho\}$ a basic system of zeroth order invariants of $\bar{Y}$, one is naturally lead to consider $\xi^{1}=x,\,\xi^{2}=t$ and by taking $\nu=z+4\,\arctan\rho$ one obtains $z^{\prime}_{xt}-\sin z^{\prime}=0$. Hence, one has the auto-Bäcklund transformation of $\mathcal{E}$
@@
provided by the invariants of the pseudosymmetry $\bar{Y}$.
@@
We consider here the short-pulse equation $\mathcal{E}=\{z_{xt}=\frac{1}{2}z^{2}z_{xx}+zz_{x}^{2}+z\}$ and show how the invariants of a nonlocal pseudosymmetry $\bar{Y}$ can be used to determine an auto-Bäcklund transformation of this equation. To this end we consider the following ZCR
$\alpha=\frac{1}{4\eta}\left(\begin{array}[]{ll}1&z_{x}\vskip 5.0pt\\ z_{x}&-1\end{array}\right)dx+\frac{1}{8\eta}\left(\begin{array}[]{ll}8\eta^{2}+z^{2}&z\left(zz_{x}-4\,\eta\right)\vskip 5.0pt\\ z\left(zz_{x}+4\,\eta\right)&-8\eta^{2}-z^{2}\end{array}\right)dt,\qquad\eta\in\mathbb{R}-\{0\}.$
@@
$\mathcal{Y}:\;\left\{\begin{array}[]{l}{\displaystyle\rho_{x}=-\frac{z_{x}}{4\eta}\rho^{2}+\frac{1}{2\eta}\,\rho+\frac{z_{x}}{4\eta}},\vskip 10.0pt\\ {\displaystyle\rho_{t}=-\frac{z\left(zz_{x}+4\,\eta\right)}{8\eta}\rho^{2}+\frac{\left(8\,\eta^{2}+z^{2}\right)}{4\eta}\rho-\frac{z\left(-zz_{x}+4\,\eta\right)}{8\eta}},\vskip 10.0pt\\ {\displaystyle z_{xt}=\frac{1}{2}z^{2}z_{xx}+zz_{x}^{2}+z,}\end{array}\right$.
@@
First, by searching for pseudosymmetries $\bar{Y}:=\left.Y\right|_{\mathcal{Y}^{(\infty)}}$ of $\mathcal{Y}$ determined by the pseudoprolongations $Y$ of vector fields of the form
@@
relatively to conservation laws $c\,\mu$, with $c\in\mathbb{R}-\{0\}$, one finds that the most general pseudosymmetry of this type with $b^{1}\neq 0$ is a constant multiple of
$\bar{Y}=-2\rho\partial_{x}+\left(\rho^{2}-1\right)\partial_{z}-\frac{1}{8\eta}\left(\rho^{2}+1\right)^{2}\partial_{\rho}+\ldots\,$,
with $c=2$. Such a pseudosymmetry of $\mathcal{Y}$ can be seen as a nonlocal pseudosymmetry of $\mathcal{E}$, with respect to the differentiable covering defined by $\mathcal{Y}$. Thus, being $\{x+8\eta/(\rho^{2}+1),\,t,z-8\eta\rho/(\rho^{2}+1)\}$ a basic system of zeroth order invariants of $\bar{Y}$, one is naturally lead to consider $\xi^{1}=x+8\eta/(\rho^{2}+1),\,\xi^{2}=t$ and by taking $\nu=-z+8\eta\rho/(\rho^{2}+1)$ one obtains $z^{\prime}_{x^{\prime}t^{\prime}}-\frac{1}{2}z^{\prime}{}^{2}z^{\prime}_{x^{\prime}x^{\prime}}-z^{\prime}z^{\prime}{}_{x^{\prime}}^{2}-z^{\prime}=0$. Hence, one has the auto-Bäcklund transformation of $\mathcal{E}$
@@
provided by the invariants of the pseudosymmetry $\bar{Y}$.
@@
We consider here Camassa-Holm equation $\mathcal{E}=\{z_{txx}=3\,z\,z_{x}-z\,z_{xxx}-2\,z_{x}\,z_{xx}+z_{t}\}$ and show how the invariants of a nonlocal pseudosymmetry $\bar{Y}$ can be used to determine an auto-Bäcklund transformation of this equation. To this end we consider the following ZCR
$\begin{array}[]{ll}\alpha={\displaystyle\frac{1}{2}}\left(\begin{array}[]{ll}0\vskip 5.0pt&-{\displaystyle\frac{1}{\eta}}\\ -z+z_{xx}-\eta&\quad 0\end{array}\right)\,dx+{\displaystyle\frac{1}{2}}\left(\begin{array}[]{ll}z_{x}\vskip 5.0pt&{\displaystyle\frac{1}{\eta}}(z-2\,\eta)\\ z^{2}-z\,z_{xx}-\eta\,z-2\eta^{2}&\quad-z_{x}\end{array}\right)dt,&{\scriptsize\eta\in\mathbb{R}-\{0\}.}\end{array}$
@@
$\mathcal{Y}:\;\left\{{\displaystyle\begin{array}[]{l}\rho_{x}={\displaystyle\frac{1}{2\eta}\rho^{2}+\frac{z_{xx}}{2}+\frac{-{\eta}^{2}-\eta\,z}{2\,\eta}},\vskip 10.0pt\\ \rho_{t}={\displaystyle-\frac{\left(-2\,\eta+z\right)}{2\,\eta}\rho^{2}-z_{x}\,\rho-\frac{z\,z_{xx}}{2}-\frac{2\,\eta^{3}+\eta^{2}z-\eta\,z^{2}}{2\,\eta}},\vskip 10.0pt\\ z_{txx}=3\,z\,z_{x}-z\,z_{xxx}-2\,z_{x}\,z_{xx}+z_{t},\end{array}}\right$. (41)
@@
In this case, if one searches for pseudosymmetries $\bar{Y}:=\left.Y\right|_{\mathcal{Y}^{(\infty)}}$ of $\mathcal{Y}$ determined by the pseudoprolongations $Y$ of (relative) vector fields of the form
@@
relatively to conservation laws $c\,\mu$, with $c\in\mathbb{R}-\{0\}$, one finds that the most general pseudosymmetry of this type with $b^{1}\neq 0$ is a constant multiple of
$\bar{Y}=-\partial_{x}+(2\rho-z_{x})\,\partial_{z}+\frac{\left(\eta^{2}-\rho^{2}\right)}{2\eta}\partial_{\rho}+\ldots\,$,
@@
$\xi^{1}=x-\ln\left(\frac{\rho-\eta}{\rho+\eta}\right),\quad\xi^{2}=t,\quad\nu^{1}=\left(\frac{\rho^{2}+\eta^{2}}{\rho^{2}-\eta^{2}}\right)z+\frac{2\,\eta\,\rho}{\rho^{2}-\eta^{2}}\,z_{x},\qquad\nu^{2}=\frac{2\,\eta\,\rho}{\rho^{2}-\eta^{2}}z+\left(\frac{\rho^{2}+\eta^{2}}{\rho^{2}-\eta^{2}}\right)\,z_{x},$
a basic system of first-order order invariants of $\bar{Y}$, one finds that in this case the system $\mathcal{Q}$ (the factorisation of $\mathcal{E}$ by $\bar{Y}$) reads
$\mathcal{Q}:\quad\left\{\begin{array}[]{l}\nu_{\xi^{1}}^{1}=\nu^{2},\vskip 10.0pt\\ \nu_{\xi^{1}\xi^{2}}^{2}=\left(-6\,\eta+3\,\nu^{1}-2\,\nu_{\xi^{1}}^{2}\right)\nu^{2}+\left(2\,\eta-\nu^{1}\right)\nu_{\xi^{1}\xi^{1}}^{2}+\nu_{\xi^{2}}^{1}.\vskip 10.0pt\end{array}\right$.
@@
one gets that $z^{\prime}_{t^{\prime}x^{\prime}x^{\prime}}-3\,z^{\prime}\,z^{\prime}_{x^{\prime}}+z^{\prime}\,z^{\prime}_{x^{\prime}x^{\prime}x^{\prime}}+2\,z^{\prime}_{x^{\prime}}\,z^{\prime}_{x^{\prime}x^{\prime}}-z^{\prime}_{t^{\prime}}=0\mod{\mathcal{E}^{(\infty)}}$. Hence, by means of (43), the invariants of $\bar{Y}$ determine an auto-Bäcklund transformation of $\mathcal{E}$. This auto-Bäcklund transformation coincides with that found in [39], by means of a different approach.